# Machine-learned trends in mirror configurations in the Large Plasma Device

### v1

@@ Abstract
This study demonstrates the efficacy of ML-based trend inference using data from the Large Plasma Device (LAPD). The LAPD is a flexible basic plasma science device with a high discharge repetition rate (0.25-1 Hz) and reproducible plasmas capable of collecting high-spatial-resolution probe measurements. A diverse dataset is collected through random sampling of LAPD operational parameters, including the magnetic field strength and profile, fueling settings, and the discharge voltage. NN ensembles with uncertainty quantification are trained to predict time-averaged ion saturation current ($I_{\text{sat}}$ — proportional to density and the square root of electron temperature) at any position within the dataset domain. Model-inferred trends, such as the effects of introducing mirrors or changing the discharge voltage, are consistent with current understanding. In addition, axial variation is optimized via comprehensive search over $I_{\text{sat}}$ predictions. Experimental validation of these optimized machine parameters demonstrate qualitative agreement, with quantitative differences attributable to Langmuir probe variation and cathode conditions. This investigation demonstrates, using ML techniques, a new way of extracting insight from experiments and novel optimization of plasmas. The code and data used in this study are made freely available.
@@ I Introduction
Many studies have used ML for profile prediction on a variety of tokamaks, particularly for real-time prediction and control. For example, NNs were used to predict electron density, temperature, and other quantities in DIII-D [1], and reservoir NNs have demonstrated the ability to quickly adapt to new scenarios or devices [2]. Temporal evolution of parameters has been successfully modeled using recurrent neural networks (RNNs)[3] for multiple devices, including the EAST[4] and KSTAR tokamaks[5, 6]. These predictions enabled training of a reinforcement learning-based controller[5, 6]. Electron temperature profiles have also been predicted using dense NNs on the J-TEXT tokamak [7].
@@
Outside of magnetized plasmas, the laser plasma community has embraced ML techniques for various applications, enumerated in a review by Dopp et al.[14]. Data-driven plasma science in general has been reviewed by Anirudh et al.[15].
@@
The goal of this study is to develop a data-driven model that can provide insight into the effect of machine parameters on plasmas produced in Large Plasma Device (LAPD) in lieu of a theoretical model. We demonstrate the capability to infer trends in a particular diagnostic signal, the time-averaged ion saturation current ($I_{\text{sat}}$), for any mirror (or anti-mirror) field geometry in a variety of machine configurations. Langmuir probes are commonly used to measure density, temperature, and potential in virtually all plasma devices in low-temperature (less than 10s of an eV) regimes. The $I_{\text{sat}}$ signal in particular is almost always used in the LAPD for calculating local plasma density.
@@ II.2 Data collection and processing of $I_{\text{sat}}$ signals
$I_{\text{sat}}$ measurements were averaged over 10 to 20 ms to exclude plasma ramp-up and fluctuations. Profile evolution is not studied to minimize computational requirements. $I_{\text{sat}}$ characteristics vary significantly between axial (z) position machine parameters. For $I_{\text{sat}}$ measurements on the same probe as a Langmuir sweep (DR2 port 26, z=863 cm), the averaging process excludes the sweep period with an additional 40 $\mu$s buffer. $I_{\text{sat}}$ measurements in DR1 that saturated either the isolation amplifier or digitizer are excluded from the dataset.
@@
These machine parameters – with the exception of gas puff duration – were randomly sampled via Latin-hypercube sampling (LHS) for 44 of the dataruns. Data were then collected with these settings. Gas puff duration was reduced for the last seven runs to 20, 10, or 5 ms. The breakdown of each setting in the dataset is given in appendix B, Table 4. The top gas puff valve was used for only the first nine dataruns of DR2 because of equipment issues.
@@ II.3 Dataset distribution and bias
Neural network inputs comprise 12 variables: source field, mirror field, midplane field, gas puff voltage, discharge voltage, gas puff duration, probe coordinates (x, y, z), probe rotation, run set identifier, and top gas puff flag. These inputs are mean-centered and normalized to the peak-to-peak value with no outliers in the dataset. The baseline models trained in section III.2 did not contain the run set identifier or top gas puff flag.
The planar measurements, constituting $\approx 64\%$ of all shots, are predominantly acquired overnight for maximal machine utilization. These longer dataruns lead to particular machine configurations being overrepresented in the dataset.
@@ III.1 Training details
where $x_{i}$ represents the input vector for the $i$th example, $y_{i}$ the target measurement, $m$ the batch size, and $f$ the NN. During training, overfitting was assessed via the validation set MSE with a traditional 80-20 train-validation random split. Unless stated otherwise, a dense neural network, 4 layers deep and 512 units wide, was trained with AdamW using a learning rate of $3\times 10^{-4}$. Leaky ReLU activations and adaptive gradient clipping[22] (cutting gradients norms above the 90th percentile of recent norms) were used to mitigate vanishing gradients and mitigate exploding gradients, respectively. The models were evaluated after training concluded at 500 epochs.
@@ III.5 Learning rate scheduling
Modifying the learning rate over time (scheduling) is known to improve model learning. The following schedules were compared: constant learning rate ($\gamma=3\times 10^{-4}$), $\gamma\propto\text{epoch}^{-1}$, $\gamma\propto\exp{(\text{epoch})}$, and $\gamma\propto\text{epoch}^{-1/2}$. The epoch is the training step divided by the number of batches in one epoch, so “epoch” in this case takes on a floating-point value. $\gamma\propto\text{epoch}^{-1}$ appears to give the best test set loss by a test MSE of $1\times 10^{-4}$, and any schedule beats a constant learning rate by $2-4\times 10^{-4}$.
@@ IV.1 $\beta$-NLL loss
[…]
@@ IV.3 Model calibration via weight decay
Increased weight decay can lead to better model calibration [27]. Model ensembles were trained with weight decay coefficients between 0 and 50 to determine the best calibrated model determined by the distribution of z-scores of the training and test sets. The results of this weight decay scan are seen in Fig. 2. Increasing the weight decay increases the test MSE and decreases its z-score standard deviation. This large standard deviation is caused by outliers. Excluding z-scores magnitudes above 10, or 4.4% of the test set, yields a standard deviation of 2.53. This long tail indicates that the distribution of predictions on the test set is not Gaussian. Nonetheless, the trend remains that increasing weight decay leads to smaller test set z-score standard deviations. However, the test set MSE increases after a weight decay of 1. This increase in test MSE implies that the model is making less accurate predictions but is better calibrated. Highly biased models are better calibrated, but come at great expense of mean prediction error. At the weight decay value of 50, the model has worse error than a linear model. Despite the attempts using weight decay, the model never becomes well-calibrated: the predicted uncertainty is always too low by a factor of 2 to 5.
@@ V.1 Checking geometrical intuition
To check this intuition, the model is given the following inputs: $B_{\text{source}}$=500 G, $B_{\text{mirror}}$=1500 G, $B_{\text{midplane}}$=500 G, discharge voltage=110 V, gas puff voltage=70 V, gas puff duration=38 ms, run set flag=DR2 and top gas puff=off. The discharge voltage and gas puffing parameters were arbitrarily chosen. The x coordinate is scanned from 0 to 30 cm, and the z coordinate from 640 to 1140 cm. This discharge is then modified by separately changing $B_{\text{source}}$ to 1500 G and $B_{\text{midplane}}$ to 750 G (M=1.5). The x profiles at the midplane (z=790 cm) of the reference M=3 prediction, source field change, and midplane field change, all scaled to cathode radius, can be seen in Fig. 4. Changing the source field to 1500 G increases the $I_{\text{sat}}$ towards the edge of the plasma, as expected. When the midplane field is increased, the $I_{\text{sat}}$ values decrease at the edge and increase at the core (x=0 cm), implying a thinner plasma column and is consistent with previously measured behavior. When only the mirror field is modified (not shown), the strongest effect on $I_{\text{sat}}$ is on or near x=0 cm, and the plasma column width does not appear to appreciably change. Strong axial $I_{\text{sat}}$ gradients are present in these predicted discharges.
@@ V.2 Directly comparing prediction to measurement
Because of data acquisition issues, only a single useful shot was collected at a nominal -45 angle (relative to the x-axis) 10 cm past the center (x=0 cm, y=0 cm) of the plasma on three probes at z-positions of 990, 767, and 607 cm (ports 22, 29, and 34, respectively). The probe drives were slightly uncentered, leading to the real coordinates of the probes to be around x $=9.75$ cm and y $=-8.4$ cm. Note that the model can predict anywhere in LAPD bounded by the training data, so off-axis measurements are not an issue. The resulting predictions using these coordinates and machine conditions can be seen in Fig. 5. The model reproduces the axial trend well, but slightly underestimates $I_{\text{sat}}$ on an absolute basis. However, given the lack of absolute $I_{\text{sat}}$ calibration and variable machine state, the agreement of the absolute $I_{\text{sat}}$ values may be coincidental. Nevertheless, the trend exhibited by this validation study match the predicted trend and increase our trust in model predictions.
Figure 5: Top: data collected at off-axis positions around x $=9.75$ cm and y $=-8.4$ cm are compared with predictions from the machine learning model at the same points in addition to two interpolating predictions. The model predicts the trend well, but underestimates $I_{\text{sat}}$ in general. The shaded orange region is the total model uncertainty ($\sigma=\sqrt{\text{Var}}$). Bottom: Measured vs predicted $I_{\text{sat}}$ values for an odd machine configuration with $B_{\text{source}}$=822 G and discharge voltage=160 V. The training data only covers discharge voltages up to 150 V The machine was also in an odd discharge state, so it’s no surprise that the predicted uncertainty bounds are very large (even greater than the test set RMSE value) and that accuracy suffers.
@@ VII Optimizing profiles
One particular issue seen in LAPD plasmas is sharp axial density and temperature gradients. Some experiments require flat gradients, such as Alfvén wave propagation studies. We explore optimizing the axial $I_{\text{sat}}$ variation as an approximation to this problem. In addition, in this case the optimization problem is used as a tool to evaluate the quality of the learned model. This is a very demanding task because the trends inferred by the model along all inputs must simultaneously be accurate. Constraints on this optimization further increase the difficulty of the problem. Success in optimization provides strong evidence that the model has inferred relevant trends in predicting $I_{\text{sat}}$. We quantify the uniformity of the axial profile by taking the standard deviation of $I_{\text{sat}}$ of 11 points along the z-axis ($x,y=0$). The required LAPD state for attaining the most uniform axial profile can be found by finding the minimum of this standard deviation with respect to the LAPD actuators (model inputs):
@@
For this optimization we use an ensemble of five $\beta$-NLL-loss models with weight decay $\lambda=0$. The $\lambda=0$ model is used because it appears to give the most useful uncertainty estimate (seen in Fig. 3). The optimal machine actuator states are found by feeding a grid of inputs into the neural network. This variance estimate is not well-calibrated: the error of the predictions on the test set falls far outside the predicted uncertainty. However, this uncertainty can be used in a relative way: when the model is predicting far outside its training range, the predicted variance is much larger. The ranges of inputs into this model are seen in Table. 2. These inputs yield 127,234,800 different machine states (times five models) which takes 151 seconds to process on an RTX 3090 ($\approx 4.2$ million forward passes per second) when implemented in a naive way. Note that gradient-based methods can be used for search because the network is differentiable everywhere but this network and parameter space is sufficiently small that a comprehensive search is computationally tractable.
@@
Since many physics studies require higher densities, we constrain the mean axial $I_{\text{sat}}$ value to be greater than 7.5 mA/mm2 (roughly $0.5\text{-}2\times 10^{13}\text{ cm}^{-3}$). The “run set flag” is set to “on” for cases to be validated (bolded in Table 3) because we wish to keep the turbopumps on to represent typical LAPD operating conditions. In addition the “top gas puff flag” was set to ‘off’ to minimize the complexity of operating the fueling system on followup dataruns and experiments. Turning the top gas puff valve on is predicted to decrease the average $I_{\text{sat}}$ by $-2$ mA/mm2 for strongly varying profiles, but otherwise the shapes are very similar. The inputs corresponding to the maximum and minimum axial variation under these constraints can be seen in columns 3 and 4 of Table 3. For contrast we also consider what machine settings would lead to the greatest axial variation. The results of both of these optimizations can be seen in Fig. 8. The optimizations yield profiles that have the largest $I_{\text{sat}}$ values closest to the cathode, which is expected.

### v2

@@ Abstract
This study demonstrates the efficacy of machine learning (ML)-based trend inference using data from the Large Plasma Device (LAPD). The LAPD is a flexible basic plasma science device with a high discharge repetition rate (0.25-1 Hz) and reproducible plasmas capable of collecting high-spatial-resolution probe measurements. A diverse dataset is collected through random sampling of LAPD operational parameters, including the magnetic field strength and profile, fueling settings, and the discharge voltage. Neural network (NN) ensembles with uncertainty quantification are trained to predict time-averaged ion saturation current ($I_{\text{sat}}$ — proportional to density and the square root of electron temperature) at any position within the dataset domain. Model-inferred trends, such as the effects of introducing mirrors or changing the discharge voltage, are consistent with current understanding. In addition, axial variation is optimized via comprehensive search over $I_{\text{sat}}$ predictions. Experimental validation of these optimized machine parameters demonstrate qualitative agreement, with quantitative differences attributable to Langmuir probe variation and cathode conditions. This investigation demonstrates, using ML techniques, a new way of extracting insight from experiments and novel optimization of plasmas. The code and data used in this study are made freely available.
@@ I Introduction
Many studies have used ML for profile prediction on a variety of tokamaks, particularly for real-time prediction and control. For example, NNs were used to predict electron density, temperature, and other quantities in DIII-D [1], and reservoir NNs have demonstrated the ability to quickly adapt to new scenarios or devices [2]. Temporal evolution of parameters has been successfully modeled using recurrent neural networks (RNNs)[3] for multiple devices, including the EAST[4] and KSTAR tokamaks[5, 6]. These predictions enabled training of a reinforcement learning-based controller[5, 6]. In addition, a decision tree-based controller was trained to maximize $\beta_{N}$ while avoiding tearing instabilities[7] on DIII-D. Electron temperature profiles have also been predicted using dense NNs on the J-TEXT tokamak [8].
@@
The use of machine learning and Bayesian inference in fusion research has been recently reviewed by Pavone et al.[15]
Outside of magnetized plasmas, the laser plasma community has embraced ML techniques for various applications, enumerated in a review by Dopp et al.[16]. Data-driven plasma science in general has been reviewed by Anirudh et al.[17] Notably, a similar quasi-random method (Sobol sequences) was used to collect a spectroscopy dataset on a plasma processing device over diverse machine settings [18]. This process is similar to what is performed in our work here, but a generative variational autoencoder was instead trained to be used as an empirical surrogate model.
@@
The goal of this study is to develop a data-driven model that can provide insight into the effect of machine parameters on plasmas produced in Large Plasma Device (LAPD) in lieu of a theoretical model. In contrast with tokamaks and other fusion devices, the LAPD is particularly well-suited for ML data collection because of its flexibility and high repetition rate. We demonstrate the capability to infer trends in a particular diagnostic signal, the time-averaged ion saturation current ($I_{\text{sat}}$), for any mirror (or anti-mirror) field geometry in a variety of machine configurations. Langmuir probes are commonly used to measure density, temperature, and potential in virtually all plasma devices in low-temperature (less than 10s of an eV) regimes. The $I_{\text{sat}}$ signal in particular is almost always used in the LAPD for calculating local plasma density.
@@ II.2 Data collection and processing of $I_{\text{sat}}$ signals
[…]
Figure 2: Gas puff timings and example $I_{\text{sat}}$ time series at three different z-axis locations from three different dataruns. Note that some discharges do not achieve steady state in $I_{\text{sat}}$.
@@
These machine parameters – with the exception of gas puff duration – were randomly sampled via Latin-hypercube sampling (LHS) for 44 of the dataruns. Data were then collected with these settings. Gas puff duration was reduced for the last seven runs to 20, 10, or 5 ms (see fig. 2 for timings relative to $I_{\text{sat}}$ signals). The breakdown of each setting in the dataset is given in appendix B, Table 4. The top gas puff valve was used for only the first nine dataruns of DR2 because of equipment issues.
@@ II.3 Dataset distribution and bias
Neural network inputs comprise 12 variables: source field, mirror field, midplane field, gas puff voltage, discharge voltage, gas puff duration, probe coordinates (x, y, z), probe rotation, run set identifier, and top gas puff flag. These variables can be interpreted as six control parameters, four probe coordinates, and two flags. These inputs are mean-centered and normalized to the peak-to-peak value with no outliers in the dataset. The baseline models trained in section III.2 did not contain the run set identifier or top gas puff flag.
Measurements over an x-y plane, constituting $\approx 64\%$ of all shots, are predominantly acquired overnight for maximal machine utilization. These longer dataruns lead to particular machine configurations being overrepresented in the dataset.
@@ III.1 Training details
where $x_{i}$ represents the input vector for the $i$th example, $y_{i}$ the target measurement, $m$ the batch size, and $f$ the NN. During training, overfitting was assessed via the validation set MSE with a traditional 80-20 train-validation random split. Unless stated otherwise, a dense neural network, 4 hidden layers deep and 256 units wide ($\approx$200,000 parameters), was trained with AdamW using a learning rate of $3\times 10^{-4}$. Leaky ReLU activations (the nonlinearities in the NN) and adaptive gradient clipping[25] (cutting gradients norms above the 90th percentile of recent norms) were used to mitigate vanishing gradients and mitigate exploding gradients, respectively. The models were evaluated after training concluded at 500 epochs.
@@ III.5 Learning rate scheduling
Modifying the learning rate over time (scheduling) is known to improve model learning. The following schedules were compared: constant learning rate ($\gamma=3\times 10^{-4}$), $\gamma\propto\text{epoch}^{-1}$, $\gamma\propto\exp{(-\text{epoch})}$, and $\gamma\propto\text{epoch}^{-1/2}$. The epoch is the training step divided by the number of batches in one epoch, so “epoch” in this case takes on a floating-point value. $\gamma\propto\text{epoch}^{-1}$ appears to give the best test set loss by a test MSE difference of $1\times 10^{-4}$, and any schedule beats a constant learning rate by a difference of $2-4\times 10^{-4}$.
@@ IV.1 $\beta$-NLL loss
This NLL-like loss assumes the prediction – the likelihood of $y$ given input $\mathbf{x}$: $p(y|\mathbf{x})$ – follows a Gaussian distribution. Treating each prediction as an independent random variable (considering each model in the ensemble is sampled from some weight distribution $\theta\sim p(\theta|\mathbf{x},y)$) and finding the mean of the random variables yields a normal distribution with mean $\mu_{*}(\mathbf{x})=\langle\mu_{i}(\mathbf{x})\rangle$ and variance $\sigma^{2}_{*}=\langle\sigma^{2}_{i}(\mathbf{x})+\mu^{2}_{i}(\mathbf{x})\rangle-\mu^{2}_{*}(\mathbf{x})$ where $\langle\rangle$ indicates an average over the ensemble.
@@ IV.2 Cross-validation MSE
All validation set MSEs fall between $1$ and $6\times 10^{-5}$, with the average training MSE falling within that range as well. These MSEs indicate that the model is able to fit the training data to a high degree of accuracy regardless of which dataruns are held out. The loss and MSE curves over training epochs can be seen in the appendix in fig. 11.
@@ IV.3 Model calibration via weight decay
Increased weight decay can lead to better model calibration [30]. Weight decay penalizes large parameter values by adding the L2 norm of model weights to the loss. Model ensembles were trained with weight decay coefficients between 0 and 50 to determine the best calibrated model determined by the distribution of z-scores of the training and test sets. The results of this weight decay scan are seen in Fig. 3. Increasing the weight decay increases the test MSE and decreases its z-score standard deviation. This large standard deviation is caused by outliers. Excluding z-scores magnitudes above 10, or 4.4% of the test set, yields a standard deviation of 2.53. This long tail indicates that the distribution of predictions on the test set is not Gaussian. Nonetheless, the trend remains that increasing weight decay leads to smaller test set z-score standard deviations. However, the test set MSE increases after a weight decay of 1. This increase in test MSE implies that the model is making less accurate predictions but is better calibrated. Highly biased models are better calibrated, but come at great expense of mean prediction error. At the weight decay value of 50, the model has worse error than a linear model. Despite the attempts using weight decay, the model never becomes well-calibrated: the predicted uncertainty is always too low by a factor of 2 to 5.
@@ V.1 Checking geometrical intuition
To check this intuition, the model is given the following inputs: $B_{\text{source}}$=500 G, $B_{\text{mirror}}$=1500 G, $B_{\text{midplane}}$=500 G, discharge voltage=110 V, gas puff voltage=70 V, gas puff duration=38 ms, run set flag=DR2 and top gas puff=off. The discharge voltage and gas puffing parameters were arbitrarily chosen. The x coordinate is scanned from 0 to 30 cm, and the z coordinate from 640 to 1140 cm. This discharge is then modified by separately changing $B_{\text{source}}$ to 1500 G and $B_{\text{midplane}}$ to 750 G (M=1.5). The x profiles at the midplane (z=790 cm) of the reference M=3 prediction, source field change, and midplane field change, all scaled to cathode radius, can be seen in Fig. 5. Changing the source field to 1500 G increases the $I_{\text{sat}}$ towards the edge of the plasma, as expected. When the midplane field is increased, the $I_{\text{sat}}$ values decrease at the edge and increase at the core (x=0 cm), implying a thinner plasma column and is consistent with previously measured behavior. When only the mirror field is modified (not shown), the strongest effect on $I_{\text{sat}}$ is on or near x=0 cm, and the plasma column width does not appear to appreciably change.
@@ V.2 Directly comparing prediction to measurement
Because of data acquisition issues, only a single useful shot was collected at a nominal -45 angle (relative to the x-axis) 10 cm past the center (x=0 cm, y=0 cm) of the plasma on three probes at z-positions of 990, 767, and 607 cm (ports 22, 29, and 34, respectively). The probe drives were slightly uncentered, leading to the real coordinates of the probes to be around x =9.75=9.75 cm and y =8.4=-8.4 cm. Note that the model can predict anywhere in LAPD bounded by the training data, so off-axis measurements are not an issue. The resulting predictions using these coordinates and machine conditions can be seen in Fig. 6. The model reproduces the axial trend well, but slightly underestimates $I_{\text{sat}}$ on an absolute basis. However, given the lack of absolute $I_{\text{sat}}$ calibration and variable machine state, the agreement of the absolute $I_{\text{sat}}$ values may be coincidental. Nevertheless, the trend exhibited by this validation study match the predicted trend and increase our trust in model predictions.
Figure 6: Top: data collected at off-axis positions around x =9.75=9.75 cm and y =8.4=-8.4 cm are compared with predictions from the machine learning model at the same points in addition to two interpolating predictions. The model predicts the trend well, but underestimates $I_{\text{sat}}$ in general. The shaded orange region is the total model uncertainty ($\sigma=\sqrt{\text{Var}}$). Bottom: Measured vs predicted $I_{\text{sat}}$ values for an odd machine configuration with $B_{\text{source}}$=822 G and discharge voltage=160 V. The training data only covers discharge voltages up to 150 V The machine was also in an odd discharge state, so it’s no surprise that the predicted uncertainty bounds are very large (even greater than the test set RMSE value) and that accuracy suffers.
@@ VII Optimizing profiles
One particular issue seen in LAPD plasmas is sharp axial density and temperature gradients. Some experiments require flat gradients, such as Alfvén wave propagation studies. We explore optimizing the axial $I_{\text{sat}}$ variation as an approximation to this problem. In addition, in this case the optimization problem is used as a tool to evaluate the quality of the learned model. This is a very demanding task because the trends inferred by the model along all inputs must simultaneously be accurate. Constraints on this optimization further increase the difficulty of the problem. Success in optimization provides strong evidence that the model has inferred relevant trends in predicting $I_{\text{sat}}$. We quantify the uniformity of the axial profile by taking the standard deviation of $I_{\text{sat}}$ of 11 points along the z-axis ($x,y=0$). The required LAPD state for attaining the most uniform axial profile can be found by finding the minimum of this standard deviation with respect to the LAPD control parameters and flags:
@@
For this optimization we use an ensemble of five $\beta$-NLL-loss models with weight decay $\lambda=0$. The $\lambda=0$ model is used because it appears to give the most useful uncertainty estimate (seen in Fig. 4). The optimal machine actuator states are found by feeding a grid of inputs into the neural network. This variance estimate is not well-calibrated: the error of the predictions on the test set falls far outside the predicted uncertainty. However, this uncertainty can be used in a relative way: when the model is predicting far outside its training range, the predicted variance is much larger. The ranges of inputs into this model are seen in Table. 2. These inputs yield 127,234,800 different machine states (times five models) which takes 151 seconds to process on an RTX 3090 ($\approx 4.2$ million forward passes per second) when implemented in a naive way. The number of forward passes can be reduced by a factor of 51 if the x value is set to 0 cm. Note that gradient-based methods can be used for search because the network is differentiable everywhere but this network and parameter space is sufficiently small that a comprehensive search is computationally tractable.
@@
Since many physics studies require higher densities, we constrain the mean axial $I_{\text{sat}}$ value to be greater than 7.5 mA/mm2 (roughly $0.5\text{-}2\times 10^{13}\text{ cm}^{-3}$). The “run set flag” is set to “on” for cases to be validated (bolded in Table 3) because we wish to keep the turbopumps on to represent typical LAPD operating conditions. In addition the “top gas puff flag” was set to ‘off’ to minimize the complexity of operating the fueling system on followup dataruns and experiments. Turning the top gas puff valve on is predicted to decrease the average $I_{\text{sat}}$ by 2-2 mA/mm2 for strongly varying profiles, but otherwise the shapes are very similar. The inputs corresponding to the maximum and minimum axial variation under these constraints can be seen in columns 3 and 4 of Table 3. For contrast we also consider what machine settings would lead to the greatest axial variation. The results of both of these optimizations can be seen in Fig. 9. The optimizations yield profiles that have the largest $I_{\text{sat}}$ values closest to the cathode, which is expected.